OpenStudy (anonymous):

Find the derivative y’(x) implicitly. x^2 y^2 + 3y = 4x

OpenStudy (anonymous):

@wio can u help?

OpenStudy (anonymous):

Sure, what part are you having trouble with?

OpenStudy (anonymous):

how do i do it? i already started but i'm stuck so i found the derivative.... f'(x) = 2x y^2 + 3 = 4 is that correct?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

ohh how do i do it then?

OpenStudy (anonymous):

For example: \[ \frac{d}{dx}3y = 3\frac{dy}{dx} \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[x^2y^2 + 3y =4x\] So now the first part is product rule. and you get \[(2xy^2 + 2x^2y \frac{ dy }{ dx }) + 3\frac{ dy }{ dx } = 4\] From there, collect the dy/dx terms on the left, and move the others to the right\[\frac{ dy }{ dx }(2x^2y + 3) = 4-2xy^2\]From there, you divide by the dy/dx term and you get\[\frac{ dy }{ dx }= \frac{ 4-2xy^2 }{ 2x^2y + 3 }\]

OpenStudy (anonymous):

\[ x^2 y^2 = \frac{d}{dx}[x^2]y^2+x^2\frac{d}{dx}y^2 = 2xy^2+x^22y\frac{dy}{dx} \]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Do you understand what is going on?

OpenStudy (anonymous):

When finding the derivative for a function of \(y\), you have to use properties of derivatives, such as chain rule.

OpenStudy (anonymous):

okay i got you